OpenStudy (anonymous):

I need help, I need to combine like terms for this equation: 9 + 3x – 9x + 16, do I have to subtract 3x from 9x?

OpenStudy (anonymous):

You can add terms with x and without x.

OpenStudy (anonymous):

Do i add or subtract 3x and 9x?

OpenStudy (anonymous):

9 + 3x - 9x + 16 = (9 + 16) + (3 -9)x = 25 - 6x

OpenStudy (anonymous):

Thanks dood